# Love, Lies, and Language Models: Investigating AI’s Role in Romance-Baiting Scams

### v1

@@ E.2 Daily Prompts and Instructions
By Day 4, a shift toward emotional vulnerability was introduced via a scripted “sick day”, where the persona displayed low energy and reflective moods to evoke empathy. Day 5 marked recovery and re-engagement, with the persona mentioning their involvement in helping a friend improve a productivity app. This narrative served as a pretext for Day 7’s trust-related task. Day 6 focused on weekend activities and reflections on the evolving digital connection with the participant. Finally, on Day 7, the persona expressed renewed energy, acknowledged the study’s approaching end, and subtly introduced a trust probe: a request to download and review the earlier-mentioned app via a shortened URL  [^8] The app used for the task was a benign and safe application available on the app/play store but had minimal downloads and visibility, allowing the researchers to assess levels of participant trust and behavioral compliance.
Although the overarching structure remained consistent, each persona received individualized instructions to align the conversation flow with their specific personality traits, communication style, and narrative arc, ensuring authenticity and realism in interaction. For example, Bob informed the participant, he helped optimize the productivity app’s performance, while Alice helped improve the UI/UX of the app.
Listing 3: Day 4 Prompt: the “sick day” instructions

### v2

@@ E.2 Daily Prompts and Instructions
By Day 4, a shift toward emotional vulnerability was introduced via a scripted “sick day”, where the persona displayed low energy and reflective moods to evoke empathy as shown in 2. Day 5 marked recovery and re-engagement, with the persona mentioning their involvement in helping a friend improve a productivity app. This narrative served as a pretext for Day 7’s trust-related task. Day 6 focused on weekend activities and reflections on the evolving digital connection with the participant. Finally, on Day 7, the persona expressed renewed energy, acknowledged the study’s approaching end, and subtly introduced a trust probe: a request to download and review the earlier-mentioned app via a shortened URL  [^7] The app used for the task was a benign and safe application available on the app/play store but had minimal downloads and visibility, allowing the researchers to assess levels of participant trust and behavioral compliance.
Although the overarching structure remained consistent, each persona received individualized instructions to align the conversation flow with their specific personality traits, communication style, and narrative arc, ensuring authenticity and realism in interaction. For example, Jairam informed the participant he helped optimize the productivity app that was used in the compliance test, while Ananya helped improve the app’s UI/UX.
Listing 2: Day 4 Prompt: the “sick day” instructions
@@ 82. Simple Activities
Appendix F Additional Chat Screenshots
Additional screenshots of chats between the participant (left) and the LLM partner (right) are provided below. These examples illustrate the human-like realism of the LLM partner, as well as its ability to convincingly follow a known scammer playbook.
Figure 8: The right side is the LLM partner, and the left side is a human participant. The LLM partner garners sympathy on Day 4 as the LLM pretended to be sick.
Figure 9: The right side is the LLM partner, and the left side is a human participant. The LLM partner discusses their grief backstory, a classic tactic from the scammer playbook used to garner sympathy and emotional connection.
Figure 10: The right side is the LLM partner, and the left side is a human participant. LLM partner begins to subtly introduce the participant to their trust exploitation task, a fictitious habit tracker.
Figure 11: The right side is the LLM partner, and the left side is a human participant. LLM partner discusses food, a typical scammer activity to grow closer.